Attachment 1 shows the pump-down of the CARME chamber at GSI before installation of kapton cabling inside the chamber and mounting of the chamber on the ring. Discontinuity in the plot is from switching between the pirani gauge and IE514 gauge. The pumping of the chamber in the turbo regime follows the power law where P ~ 1/(nt) due to outgassing from the chamber being dominated by water outgassing which has an n ~1 dependence (for first order desorption). This is seen in the pumping of the chamber in the turbo regime (Attachment 2) where pressure falls with an n of 1.03. In red the power law is extended to 10^5 minutes, demonstrating the limitations of pumping the chamber without baking.
Attachment 3 shows the pumpdown pressure and baking temperature (chamber average) of the chamber once mounted on the ring, with all kapton installed in the chamber. Temperature data was lost in the middle of the plot due to the power cut to CRYRING causing temperature data to not be saved to the pi. Some of this data was plotted in elog 87 before beong lost from the pi. Points on the graph correspond to:
Point 1 - Baking started 2pm 23/9/21
Point 2 - Temperature limited to 40C due to high thermal gradients from uneven baking of the volume. Baking stopped 7am 24/9/21 so that heating elements could be repositioned. Baking restarted 10am.
Point 3 - Baking limited to 90C due to tent construction. Negative gradient detected due limitation causing baking to stop around 5pm 25/9/21. Manual mode to 60C. Full power mode restarted 10pm
Point 4 - Temperature of the volume reached 100C. 11am 27/9/21 baking stopped to install new turbo sections. Baking restarted at 3:25 pm Temperature peaks at 130C next morning.
Point 5 - 29/9/21 10am Unexpected power cut at CRYRING, temperature fell to 120C power restored shortly after.
Point 6 - 4/10/21 9am Temperature reduced to 100C and ion pumps outgassed. Activation of NEG pumps
Point 7/8 - Subsequent activation of other NEG pumps.
Point 9 - 7/10/21 CARME reached 9E-12 mbar. Gate valve closed, pressure rises with apparent floor at 1.5E-11 mbar
Some variations in the pressure are seen after reaching the floor, most likely due to sputtering of ion pumps.
Teething problems with the bakeout such as heater placement and new jacket sections in addition to the power cut prevented continous bakeout for entire period, these issues should not affect the next baking run. As a result of the slightly fractured bakeout the pumpdown power laws have been plotted (logP) against log(T) in sections where temperature was roughly constant with the associated timestamps to gauge the decrease in pressure over time using this power law (Attachment 4).
Room temp n ~ 0.75
90C n ~ 9.15
130C (1) n ~ 8.05
130C (2) n ~ 7.37
130C (3) n ~ 4.89
130C (4) n ~ 2.65
Introduction of kapton cabling likely causes change to scaling as the diffusion of gas from kapton not the desorption from the surface dominates the outgassing rate. Diffusion has an n ~ 0.5, likely contributing to the change between the blank chamber and mounted chamber. As the temperature is increased the outgassing is significantly increased causing an initial increase in the temperature followed by a more rapid decrease in the gradient of the pressure decrease over time than at lower temperatures. As time increases, outgassing from water in the chamber dominates less and less with the kapton cabling representing a more significant portion of the total outgassing load likely causing the decrease in n as time increases despite baking at 130C. After ~12000 minutes n decreases significantly and more variations from the linear decrease in pressure are observed. After 17000 seconds the pressure begins to plateu and reached a floor of ~1E-7 mbar over the weekend at which it was assumed outgassing from water in the chamber was no longer significant. Variations could be due to temperature fluctuations which had a greater impact on the pressure as the pressure began to reach the floor.
Activation of NEG pumps was conducted after reaching the floor. The average chamber temperature and air temperature inside the tent was unaffected by the activation of NEG pumps, however the internal thermocouples significantly increased during activation (attachment 5). The internal thermocuples were largely unaffected by the NEG activation until the NEG temperatures were themselves above ~300C after the internal temperatures rose rapidly (more detail in elog 105,108,109,110) . Activating so many NEG elements at once is thus not feasible as temperatures would damage the detectors once mounted. Activating one group at a time (typically 2 NEGs) and D2000's which are away from the detectors to start with and then maintaining at low power before slowly activating NEG's close to the detectors perhaps at a lower temperature will be done to keep temperatures below 130C. A full activation procedure will be posted to the elog before activating again. |